OpenStudy (anonymous):

find the domain and range of each function. (1.) t(x)=cos4x (2.) H(G(x))

OpenStudy (anonymous):

1) domain of \[t(x)=\cos(4x)\] is \[(-\infty, \infty)\] and its range is \[(-1, 1)\] 2) H(G(x)) is composite function, functions were not given, but domain of H is range of G(x)

OpenStudy (anonymous):

Trig functions are "repeating" so they can go on forever, therefore it's domain is from negative infinity to positive infinity. The "normal" range of cos x functions is from [-1,1].. Therefore, it actually includes 1..

OpenStudy (anonymous):

yes, range including end points -1, and 1